# ARCANE: Adaptive Routing with Caching and Aware Network Exploration

### v4

@@ 4.4.2. Asymmetric Network Conditions
We evaluate the performance of ARCANE under asymmetric network conditions. We connect 16 endpoints through two T0 switches (8 endpoints each) with a total of 4 links to a pair of T1 switches. To demonstrate the adaptive load balancing capabilities of ARCANE, we change the link speed of one T0-T1 link from 400 Gbps to 200 Gbps, creating asymmetry in the network. Figure 9(b) shows the per-flow goodput as observed by the application while Figure 10(a) the FCT distribution. OPS sends packets across all paths (including those crossing the 200 Gbps link) with equal probability and is ultimately capped by the slower 200 Gbps path. The ECN marking on the 200 Gbps path causes the CC algorithm to throttle all flows and eventually match the capacity of that single slower link, thus leading to underutilization of the remaining 400 Gbps links (that are running at 50% utilization).
ARCANE can gracefully adapt in such a scenario as the cached entropies will reflect the network asymmetry and result in a path distribution that is skewed to tailor to the relative capacity of the available paths. In this example, ARCANE can reach high utilization with average per-flow goodput within 5% of the ideal fair-share target.
@@
Figure 9. ARCANE-FPGA impact on goodput.
@@
Figure 10. ARCANE-FPGA impact on FCT and packet drops
@@ 4.4.3. Network Failures
We also evaluate the performance of ARCANE in the presence of network failures. To demonstrate the robustness and resilience of ARCANE in the context of link failures, Figure 10(b) shows total packet drops (average across five runs each) observed in a large-scale 128 endpoint run (endpoints split across 2 T0s connected through 8 T1s) where we abruptly bring down a T0-T1 link during the experiment. While the network is trying to recover from the impact of this event (which in our environment can take in the order of 100s of milliseconds), OPS continues sending packets across all paths (including those affected by the link that went down). The freezing capability of ARCANE can quickly adapt to such events (within the order of an RTO) and avoid sending packets down the affected paths as the entropy cache is replenished from packets traversing healthy paths.
4.5. ARCANE Applicability
In this section, we briefly evaluate ARCANE, in simulations, under different scenarios by changing the ACK coalescing ratio, the EVS size and, the underlying CC algorithm. Generally, we believe that ARCANE can work well under many different circumstances, topologies and workloads.
@@ 4.5.1. ACK Coalescing
We have primarily evaluated ARCANE without ACK coalescing, as this configuration allows ARCANE to operate with the most up to date data. However, some transport protocols permit ACK coalescing, where the receiver sends an ACK packet only after receiving $n$ data packets from the sender. In theory, the coalesced ACK packet could return all previous non-ECN marked entropies in its header, but we focus on the worst case scenario for ARCANE where the ACK only carries the entropy of the packet that triggered it.
@@
We run a real 8 MiB permutation workload simulating different ACK coalescing ratios. As we can see in the left part of Figure 11, ARCANE with 2:1, 4:1 and 8:1 coalescing ratios does significantly better than OPS, while with 16:1 it starts losing its advantage. However, it should be noted that in case of asymmetries or failures (right Figure 11), ARCANE does much better due to its various improvements over OPS even at 16:1 ratio. We confirm these results theoretically in Appendix D.1.
@@ 4.5.2. EVS Size
In Appendix B, we prove that OPS routing, done through EVs and ECMP hashing, requires a large EVS to work correctly and reduce collisions. Moreover, this requirement also grows with the number of output ports in a switch as demonstrated in Section 5.1. On the other hand, ARCANE, due to its adaptive nature, can drastically reduce the EVS size and still work well. We show this in the left plot of Figure 12 where we compare, in a real scenario, OPS and ARCANE when using 32, 256, and 64K EVs. ARCANE works equally well with 256 and 64K EVs and is only 8% slower with 32 EVs. On the other hand, OPS is 21% and 64% slower with 256 and 32 EVs when compared to 64K. This confirms that ARCANE could potentially work well even just with 1 byte for the EVS. We note that while OPS could be implemented without using EVS—such as through round-robin selection or by making random choices directly at the switch—these approaches introduce additional challenges (Section 5.2).
@@ 4.5.3. Different CC Algorithms
ARCANE has been designed to work with any CC algorithm as long as there is no over-reaction to out-of-order packets and ECN support. In this paper, we have, so far, evaluated ARCANE working alongside a tuned version of DCTCP. However, as we will see, ARCANE can work well even with other algorithms. Moreover, we envision also a version of ARCANE that could work just with delay if ECN is not supported but we do not go into details here. For example, in the right plot of Figure 12, we run a 8 MiB permutation workload without failure for DCTCP, EQDS and a proprietary CC algorithm. ARCANE can help all of these CC algorithms when compared to OPS.
Figure 12. Performance with different EVs sizes and CCs during a 8MiB permutation.
@@ 5. Theoretical Verification
To support our experimental findings, we present a theoretical first principles analysis of OPS and demonstrate how it can lead to arbitrarily long queues. To address this issue, we develop a new recycled balls-into-bins model and prove its local convergence. This model serves as the theoretical intuition behind the ARCANE protocol (Section 3).
@@ 5.1. Recycled Balls-into-Bins Model
At maximum injection rate, OPS suffers from severe load-imbalance that eventually leads to queues of arbitrary length building up. We explain this behavior with an infinite batched balls-into-bins model (Berenbrink et al., 2018; Becchetti et al., 2019; Los and Sauerwald, 2023). In contrast, we show that recycling good paths such as in ARCANE leads to a convergent behavior and logarithmically-bounded queues.
@@
For $n\geq 16$, $\tau\geq 4\ln n$, $b\geq 2.4\ln n$, recycled balls-into-bins converges in $O(n\log n)$ steps. Every bin has $O(\log n)$ elements throughout, with probability $1-o(1)$.
@@ 5.2. Limitations and Alternatives
While the provided model performs well in an idealized setting with maximal injection rates, in a real network scenario, CC algorithms would reduce the sending rate if a large queue were to build up. However, such activations of the CC would result in a slowdown of the sending rate and, hence, an increase of the completion time of a workload as we can see in Section 4.3.1. Moreover, we note that while real scenarios would have more subtle interactions not modelled here (like using distinct $KMin$ and $KMax$), the key results hold and are consistent with our evaluations in Section 4. Finally, we note that here we provide a proof only about the local convergence of ARCANE and that other models such as static path assignments and round-robin across the ports would also work well and potentially provide no queueing. However, such approaches are not viable in reality for a series of reasons: they cannot react well to partial or total failures, they would struggle with multi-tier topologies and they require detailed knowledge of the underlying workloads (Gangidi et al., 2024).
@@ 6. Related Work
The literature for load balancing designs is vast, however most of it focuses on TCP-based deployments where out-of-order packets are not desirable (Benson et al., 2011; Zhang et al., 2017; Al-Fares et al., 2010; Vanini et al., 2017; Dixit et al., 2013). In contrast, ARCANE aligns with a new transport paradigm (UltraEthernet (Consortium, 2024), Adaptive RoCE by NVIDIA (NVIDIA, 2024), Falcon (goo, 2023)) where out-of-order packets help to leverage the capacity of multi-path networks.
@@
Load balancers with per-packet granularity, such as OPS (Dixit et al., 2013) or DRILL (Ghorbani et al., 2017), help drastically reduce ECMP collisions, but are still oblivious to asymmetries (for OPS) or require switch support (for DRILL). As demonstrated in Section 5 OPS can suffer even in symmetric cases. MPRDMA (Lu et al., 2018) also uses ECN for load balancing like ARCANE, however it requires probing and ACK clocking and does not offer caching of entropies. Hermes (Zhang et al., 2017) combines both ECN and delay but it works best with TCP-like protocols and has many parameters with complex tuning.
@@
Proteus (Hu et al., 2023) focuses on optimizing the load balancing of lossless PFC (Priority Flow Control) networks while ARCANE focuses on lossy networks.
@@ 7. Conclusion
We presented ARCANE, a simple, lightweight, yet highly effective load-balancing mechanism designed to meet the constraints of next-generation datacenter networks tailored for AI workloads. As demonstrated in our extensive evaluations, conducted through both simulations and FPGA hardware, ARCANE’ adaptive entropy caching enhances end-to-end performance across multiple critical metrics, including average flow completion time, runtime, and packet loss. ARCANE outperforms ECMP and OPS by up to 6x and 1.25x in symmetric networks, and by up to 4.5x and 2x in asymmetric networks, outperforming OPS by as much as 100x during short-term transient link failures while reducing packet drops by over 70x. Moreover, ARCANE consistently beats or matches other state-of-the-art algorithms in all scenarios evaluated. We also showed how ARCANE can work well under various network configurations demonstrating its flexibility to adapt to different scenarios while remaining light-weight on NIC memory requiring only 25 bytes of state per-connection.
Appendix A Freezing Mode in ARCANE
Ideally, ARCANE should enter freezing mode only upon detecting a network failure. To achieve this, we employ two strategies:
@@ Appendix A Freezing Mode in ARCANE
In our paper, we focused on scenarios where packet trimming was not supported. However, ARCANE performs optimally when packet trimming is available, benefiting from both an enhanced loss detection algorithm and a more responsive CC loop.
Regardless of the employed strategy, ARCANE maintains high performance even if it inadvertently enters freezing mode without an actual network failure. This is because entering freezing mode effectively reduces the EVS size of ARCANE. As demonstrated in Section 4.5.2, ARCANE remains effective with as few as 32 EVs. For instance, we tested ARCANE with the 16 MiB tornado workload, running it first under normal conditions and then with forced freezing mode activated after $150\;\mu\text{s}$. The results showed only a 1% increase in completion time for ARCANE with forced freezing mode, with the 99th percentile RTT remaining very close, as illustrated in Figure 15.
As an extension of ARCANE, probing can be incorporated to enhance failure detection precision. However, we chose not to integrate this into ARCANE at this stage to maintain clarity and focus on the core framework. We also do not discuss fast loss-recovery mechanisms as they are anyway orthogonal to ARCANE behaviour and could be used to further improve its performance.
Figure 15. Studying ARCANE while forcing freezing mode without failures.
@@ Appendix B Entropy Values Set Analysis
We first show theoretically that the EVS size is very important for OPS and later that ARCANE is much more flexible and can work well even with small EVS.
@@
Appendix C ARCANE Single Switch Convergence Proof
@@
Second Phase At the beginning of the second phase, we have at most $2n\ln n+n$ balls in total and all bins are nonempty. We define a potential function to measure the drift towards a completely load-balanced state. Let $X^{i}_{t}$ be the number of balls in bin $i$ at time step $t$ and let $Y^{i}_{t}=\max(0,X^{i}_{t}-\tau)$. Our potential is $Y_{t}=\sum_{i}Y^{i}_{t}$. Clearly, when the potential is 00, all bins have at most $\tau$ elements. We need to bound the drift $Y_{t+1}-Y_{t}$.
@@
$E[Y^{i}_{t+1}-Y^{i}_{t}|X^{i}_{t}=x,K_{t}=k]=\begin{cases}\frac{k}{n}-1&\text{if }x>\tau,\\ \frac{k}{n}&\text{if }x=\tau\\ 0&\text{if }x\leq\tau-k,\\ \end{cases}$
@@
$E[Y_{t+1}-Y_{t}|K_{t}=k]\leq\begin{cases}\frac{1}{\ln n}-\frac{k}{4}&\text{if }k>\ln n,\\ -\frac{1}{4}&\text{else.}\\ \end{cases}$
@@
Note that for $n\geq 16$, $\frac{e^{5/4}}{n}+\frac{1}{\ln n}-\frac{1+\ln n}{4}\geq-\frac{1}{8}$ and $\frac{e^{5/4}}{n}-\frac{1}{4}\geq-\frac{1}{32}$. Hence, the expected drift is at most $-\frac{1}{32}$. We conclude by an additive drift theorem (Lengler, 2020; He and Yao, 2004) that the expected time until all bins are below the threshold is at most $E[32\cdot Y_{0}]=O(n\ln n)$. If all balls have unique colors at the end, the process converges then. This happens with probability $1-O(\frac{\ln n}{n})$. So with high probability, a single such $Y_{t}=0$ event suffices and the overall expectation is $O(n\ln n)$.
@@
Let ${B}$ be binomially distributed with parameters $\frac{1}{n}$ and $k\leq n$. Then, for any $x\geq 16$:
@@
$\leq e^{-\frac{5}{4}\delta}$ using $\text{using }\delta\geq 15$
@@
Let ${B}$ be binomially distributed with parameters $\frac{1}{n}$ and $k\leq n/2$. Then, for any $x\geq 7/2$:
@@
$\leq e^{-\frac{\delta}{\frac{1}{\delta}+\frac{2}{3}}}$ using $\text{using }E[B]\leq\frac{1}{2}$
$\leq e^{-\delta}$ using $\text{using }\delta\geq 3$

### v5

@@ 4.4.2. Asymmetric Network Conditions
We evaluate the performance of REPS under asymmetric network conditions. We connect 16 endpoints through two T0 switches (8 endpoints each) with a total of 4 links to a pair of T1 switches. To demonstrate the adaptive load balancing capabilities of REPS, we change the link speed of one T0-T1 link from 400 Gbps to 200 Gbps, creating asymmetry in the network. Figure 9(b) shows the per-flow goodput as observed by the application while Figure 10(a) the FCT distribution. OPS sends packets across all paths (including those crossing the 200 Gbps link) with equal probability and is ultimately capped by the slower 200 Gbps path. The ECN marking on the 200 Gbps path causes the CC algorithm to throttle all flows and eventually match the capacity of that single slower link, thus leading to underutilization of the remaining 400 Gbps links (that are running at 50% utilization).
REPS can gracefully adapt in such a scenario as the cached entropies will reflect the network asymmetry and result in a path distribution that is skewed to tailor to the relative capacity of the available paths. In this example, REPS can reach high utilization with average per-flow goodput within 5% of the ideal fair-share target.
@@
Figure 9. REPS-FPGA impact on goodput.
@@
Figure 10. REPS-FPGA impact on FCT and packet drops
@@ 4.4.3. Network Failures
We also evaluate the performance of REPS in the presence of network failures. To demonstrate the robustness and resilience of REPS in the context of link failures, Figure 10(b) shows total packet drops (average across five runs each) observed in a large-scale 128 endpoint run (endpoints split across 2 T0s connected through 8 T1s) where we abruptly bring down a T0-T1 link during the experiment. While the network is trying to recover from the impact of this event (which in our environment can take in the order of 100s of milliseconds), OPS continues sending packets across all paths (including those affected by the link that went down). The freezing capability of REPS can quickly adapt to such events (within the order of an RTO) and avoid sending packets down the affected paths as the entropy cache is replenished from packets traversing healthy paths.
4.5. REPS Applicability
In this section, we briefly evaluate REPS, in simulations, under different scenarios by changing the ACK coalescing ratio, the EVS size and, the underlying CC algorithm. Generally, we believe that REPS can work well under many different circumstances, topologies and workloads.
@@ 4.5.1. ACK Coalescing
We have primarily evaluated REPS without ACK coalescing, as this configuration allows REPS to operate with the most up to date data. However, some transport protocols permit ACK coalescing, where the receiver sends an ACK packet only after receiving $n$ data packets from the sender. In theory, the coalesced ACK packet could return all previous non-ECN marked entropies in its header, but we focus on the worst case scenario for REPS where the ACK only carries the entropy of the packet that triggered it.
@@
We run a real 8 MiB permutation workload simulating different ACK coalescing ratios. As we can see in the left part of Figure 11, REPS with 2:1, 4:1 and 8:1 coalescing ratios does significantly better than OPS, while with 16:1 it starts losing its advantage. However, it should be noted that in case of asymmetries or failures (right Figure 11), REPS does much better due to its various improvements over OPS even at 16:1 ratio. We confirm these results theoretically in Appendix D.1.
@@ 4.5.2. EVS Size
In Appendix B, we prove that OPS routing, done through EVs and ECMP hashing, requires a large EVS to work correctly and reduce collisions. Moreover, this requirement also grows with the number of output ports in a switch as demonstrated in Section 5.1. On the other hand, REPS, due to its adaptive nature, can drastically reduce the EVS size and still work well. We show this in the left plot of Figure 12 where we compare, in a real scenario, OPS and REPS when using 32, 256, and 64K EVs. REPS works equally well with 256 and 64K EVs and is only 8% slower with 32 EVs. On the other hand, OPS is 21% and 64% slower with 256 and 32 EVs when compared to 64K. This confirms that REPS could potentially work well even just with 1 byte for the EVS. We note that while OPS could be implemented without using EVS—such as through round-robin selection or by making random choices directly at the switch—these approaches introduce additional challenges (Section 5.2).
@@ 4.5.3. Different CC Algorithms
REPS has been designed to work with any CC algorithm as long as there is no over-reaction to out-of-order packets and ECN support. In this paper, we have, so far, evaluated REPS working alongside a tuned version of DCTCP. However, as we will see, REPS can work well even with other algorithms. Moreover, we envision also a version of REPS that could work just with delay if ECN is not supported but we do not go into details here. For example, in the right plot of Figure 12, we run a 8 MiB permutation workload without failure for DCTCP, EQDS and a proprietary CC algorithm. REPS can help all of these CC algorithms when compared to OPS.
Figure 12. Performance with different EVs sizes and $C$s during a 8MiB permutation.
@@ 5. Theoretical Verification
To support our experimental findings, we present a theoretical first principles analysis of OPS and demonstrate how it can lead to arbitrarily long queues. To address this issue, we develop a new recycled balls-into-bins model and prove its local convergence. This model serves as the theoretical intuition behind the REPS protocol (Section 3).
@@ 5.1. Recycled Balls-into-Bins Model
At maximum injection rate, OPS suffers from severe load-imbalance that eventually leads to queues of arbitrary length building up. We explain this behavior with an infinite batched balls-into-bins model (Berenbrink et al., 2018; Becchetti et al., 2019; Los and Sauerwald, 2023). In contrast, we show that recycling good paths such as in REPS leads to a convergent behavior and logarithmically-bounded queues.
@@
For $n\geq 16$, $\tau\geq 4\ln n$, $b\geq 2.4\ln n$, recycled balls-into-bins converges in $O(n\log n)$ steps. Every bin has $O(\log n)$ elements throughout, with probability 1o(1)1-o(1).
@@ 5.2. Limitations and Alternatives
While the provided model performs well in an idealized setting with maximal injection rates, in a real network scenario, CC algorithms would reduce the sending rate if a large queue were to build up. However, such activations of the CC would result in a slowdown of the sending rate and, hence, an increase of the completion time of a workload as we can see in Section 4.3.1. Moreover, we note that while real scenarios would have more subtle interactions not modelled here (like using distinct $KMin$ and $KMax$), the key results hold and are consistent with our evaluations in Section 4. Finally, we note that here we provide a proof only about the local convergence of REPS and that other models such as static path assignments and round-robin across the ports would also work well and potentially provide no queueing. However, such approaches are not viable in reality for a series of reasons: they cannot react well to partial or total failures, they would struggle with multi-tier topologies and they require detailed knowledge of the underlying workloads (Gangidi et al., 2024).
@@ 6. Related Work
The literature for load balancing designs is vast, however most of it focuses on TCP-based deployments where out-of-order packets are not desirable (Benson et al., 2011; Zhang et al., 2017; Al-Fares et al., 2010; Vanini et al., 2017; Dixit et al., 2013). In contrast, REPS aligns with a new transport paradigm (UltraEthernet (Consortium, 2024), Adaptive RoCE by NVIDIA (NVIDIA, 2024), Falcon (goo, 2023)) where out-of-order packets help to leverage the capacity of multi-path networks.
@@
Load balancers with per-packet granularity, such as OPS (Dixit et al., 2013) or DRILL (Ghorbani et al., 2017), help drastically reduce ECMP collisions, but are still oblivious to asymmetries (for OPS) or require switch support (for DRILL). As demonstrated in Section 5 OPS can suffer even in symmetric cases. MPRDMA (Lu et al., 2018) also uses ECN for load balancing like REPS, however it requires probing and ACK clocking and does not offer caching of entropies. Hermes (Zhang et al., 2017) combines both ECN and delay but it works best with TCP-like protocols and has many parameters with complex tuning.
@@
Proteus (Hu et al., 2023) focuses on optimizing the load balancing of lossless PFC (Priority Flow Control) networks while REPS focuses on lossy networks.
@@ 7. Conclusion
We presented REPS, a simple, lightweight, yet highly effective load-balancing mechanism designed to meet the constraints of next-generation datacenter networks tailored for AI workloads. As demonstrated in our extensive evaluations, conducted through both simulations and FPGA hardware, REPS’ adaptive entropy caching enhances end-to-end performance across multiple critical metrics, including average flow completion time, runtime, and packet loss. REPS outperforms ECMP and OPS by up to 6x and 1.25x in symmetric networks, and by up to 4.5x and 2x in asymmetric networks, outperforming OPS by as much as 100x during short-term transient link failures while reducing packet drops by over 70x. Moreover, REPS consistently beats or matches other state-of-the-art algorithms in all scenarios evaluated. We also showed how REPS can work well under various network configurations demonstrating its flexibility to adapt to different scenarios while remaining light-weight on NIC memory requiring only 25 bytes of state per-connection.
Acknowledgements.
We thank the Swiss National Supercomputing Center (CSCS) for providing the computational resources used in this work. The authors used ChatGPT to only assist with editing and quality control throughout this manuscript (ideas and content remain original).
Appendix A Freezing Mode in REPS
Ideally, REPS should enter freezing mode only upon detecting a network failure. To achieve this, we employ two strategies:
@@ Appendix A Freezing Mode in REPS
In our paper, we focused on scenarios where packet trimming was not supported. However, REPS performs optimally when packet trimming is available, benefiting from both an enhanced loss detection algorithm and a more responsive CC loop.
Regardless of the employed strategy, REPS maintains high performance even if it inadvertently enters freezing mode without an actual network failure. This is because entering freezing mode effectively reduces the EVS size of REPS. As demonstrated in Section 4.5.2, REPS remains effective with as few as 32 EVs. For instance, we tested REPS with the 16 MiB tornado workload, running it first under normal conditions and then with forced freezing mode activated after $150\;\mu\text{s}$. The results showed only a 1% increase in completion time for REPS with forced freezing mode, with the 99th percentile RTT remaining very close, as illustrated in Figure 15.
As an extension of REPS, probing can be incorporated to enhance failure detection precision. However, we chose not to integrate this into REPS at this stage to maintain clarity and focus on the core framework. We also do not discuss fast loss-recovery mechanisms as they are anyway orthogonal to REPS behaviour and could be used to further improve its performance.
Figure 15. Studying REPS while forcing freezing mode without failures.
@@ Appendix B Entropy Values Set Analysis
We first show theoretically that the EVS size is very important for OPS and later that REPS is much more flexible and can work well even with small EVS.
@@
Appendix C REPS Single Switch Convergence Proof
@@
Second Phase At the beginning of the second phase, we have at most $2n\ln n+n$ balls in total and all bins are nonempty. We define a potential function to measure the drift towards a completely load-balanced state. Let $X^{i}_{t}$ be the number of balls in bin $i$ at time step $t$ and let $Y^{i}_{t}=\max(0,X^{i}_{t}-\tau)$. Our potential is $Y_{t}=\sum_{i}Y^{i}_{t}$. Clearly, when the potential is 0, all bins have at most $\tau$ elements. We need to bound the drift $Y_{t+1}-Y_{t}$.
@@
[…]
@@
[…]
@@
Note that for $n\geq 16$, $\frac{e^{5/4}}{n}+\frac{1}{\ln n}-\frac{1+\ln n}{4}\geq-\frac{1}{8}$ and $\frac{e^{5/4}}{n}-\frac{1}{4}\geq-\frac{1}{32}$. Hence, the expected drift is at most $\frac{1}{32}$. We conclude by an additive drift theorem (Lengler, 2020; He and Yao, 2004) that the expected time until all bins are below the threshold is at most $E[32\cdot Y_{0}]=O(n\ln n)$. If all balls have unique colors at the end, the process converges then. This happens with probability $1-O(\frac{\ln n}{n})$. So with high probability, a single such $Y_{t}=0$ event suffices and the overall expectation is $O(n\ln n)$.
@@
Let B{B} be binomially distributed with parameters $\frac{1}{n}$ and $k\leq n$. Then, for any $x\geq 16$:
@@
$\leq e^{-\frac{5}{4}\delta}$ $\text{using }\delta\geq 15$
@@
Let B{B} be binomially distributed with parameters $\frac{1}{n}$ and $k\leq n/2$. Then, for any $x\geq 7/2$:
@@
$\leq e^{-\frac{\delta}{\frac{1}{\delta}+\frac{2}{3}}}$ $\text{using }E[B]\leq\frac{1}{2}$
$\leq e^{-\delta}$ $\text{using }\delta\geq 3$